# OneRef: Unified One-tower Expression Grounding and Segmentation with Mask Referring Modeling

### v1

@@ 3.4 Referring-based grounding and segmentation transfer
OneRef-L (ours) NeurIPS’24 BEiT3-L / BEiT3-L 80.09 82.19 77.51 75.17 79.38 70.17 73.18 72.76
@@ 4.1 Experimental setups
Datasets and evaluation metrics. Our method is validated in the REC, RES, and PG tasks with five widely used datasets, namely three REC/RES datasets (RefCOCO/+/g [101, 62]), as well as two PG datasets (ReferItGame [34] and Flickr30k Entities [68]). In PG, the query pertains to a specific phrase, while in REC and RES, the query refers to a reference expression. The text of RefCOCO+/g exhibits greater length and complexity in comparison to that of RefCOCO. In REC/PG, we follow previous works [14, 96] that employs Intersection-over-Union (IoU) as the evaluation metric, i.e., a prediction is deemed accurate only when its IoU exceeds or equals 0.5. We compute the prediction accuracy for each dataset as a performance indicator. While in RES, we follow previous works [79, 35] that employs mean IoU (mIoU) for each dataset as the indicator. The detailed statistics information regarding these five datasets are provided in the Appendix B.
@@ Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62036012, U23A20387, 62322212, 62072455, in part by Pengcheng Laboratory Research Project under Grant PCL2023A08, and also in part by National Science and Technology Major Project under Grant 2021ZD0112200.
@@ E.1 The results on ReferIt and Flickr30k dataset under mixup pre-training setting
E.2 Computational costs analysis compared with SoTA methods
@@
E.3 Complete ablation study of MRefM on single-dataset fine-tuning and mixup pretrain settings
@@
E.4 Ablation study of the mask ratio in referring-aware dynamic masking strategy

### v2

@@ 3.4 Referring-based grounding and segmentation transfer
OneRef-L (ours) NeurIPS’24 BEiT3-L / BEiT3-L 80.09 82.19 77.51 75.17 79.38 70.17 73.18 73.76
@@ 4.1 Experimental setups
Datasets and evaluation metrics. Our method is validated in the REC, RES, and PG tasks with five widely used datasets, namely three REC/RES datasets (RefCOCO/+/g [101, 62]), as well as two PG datasets (ReferItGame [34] and Flickr30k Entities [68]). In PG, the query pertains to a specific phrase, while in REC and RES, the query refers to a reference expression. The text of RefCOCO+/g exhibits greater length and complexity in comparison to that of RefCOCO. In REC/PG, we follow previous works [14, 96] that employs Intersection-over-Union (IoU) as the evaluation metric, i.e., a prediction is deemed accurate only when its IoU exceeds or equals 0.5. We compute the prediction accuracy for each dataset as a performance indicator. While in RES, we follow previous works [79, 35] that employs mean IoU (mIoU) and overall IoU (oIoU) for each dataset as the indicators. The detailed statistics information regarding these five datasets are provided in the Appendix B.
@@ Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 62036012, U23A20387, 62322212, 62072455, in part by Pengcheng Laboratory Research Project under Grant PCL2023A08, in part by National Science and Technology Major Project under Grant 2021ZD0112200, in part by Alibaba Innovative Research Program, and also in part by CAS Project for Young Scientists in Basic Research (YSBR-116).
@@ Appendix E Extra experimental results
Table 13: Comparison with latest SoTA methods (oIoU metric) on the three datasets for RES task with both single-dataset fine-tuning setting and dataset-mixed intermediate pre-training setting. $\dagger$ indicates RefC has been used during pre-training.
Backbone pretrain data val testA testB val testA testB val test
LAVT [94] CVPR’22 Swin-B / BERT-B 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
RISCLIP-B [35] NAACL’24 CLIP-B / CLIP-B 73.57 76.46 69.76 65.53 70.61 55.19 64.10 65.09
RISCLIP-L [35] NAACL’24 CLIP-L / CLIP-L 76.92 80.99 73.04 71.24 76.99 61.56 67.96 68.71
OneRef-B (ours) NeurIPS’24 BEiT3-B / BEiT3-B 77.55 80.96 73.53 70.82 74.53 64.06 70.68 70.61
OneRef-L (ours) NeurIPS’24 BEiT3-L / BEiT3-L 80.48 82.78 78.27 74.25 78.41 69.85 74.91 77.36
Dataset-mixed intermediate pre-training setting:
HIPIE [88] NeurIPS’23 RN50,CLIP / BERT-B RefC,O365,PACO 78.30 66.20 69.80
UNINEXT [93] CVPR’23 RN50 / BERT-B RefC,O365,SOT 77.90 79.68 75.77 66.20 71.22 59.01 70.04 70.52
PolyFormer-B[53] CVPR’23 Swin-B / BERT-B RefC 74.82 76.64 71.06 67.64 72.89 59.33 67.76 69.05
PolyFormer-L[53] CVPR’23 Swin-L / BERT-B RefC 75.96 78.29 73.25 69.33 74.56 61.87 69.20 70.19
OneRef-B (ours) NeurIPS’24 BEiT3-B / BEiT3-B RefC 81.06 83.05 77.80 72.24 77.32 67.08 75.14 77.21
OneRef-L (ours) NeurIPS’24 BEiT3-L / BEiT3-L RefC 81.95 83.62 79.57 74.55 79.54 70.65 76.53 79.18
@@ E.2 The results for RES task under oIoU metric
E.3 Computational costs analysis compared with SoTA methods
@@
E.4 Complete ablation study of MRefM on single-dataset fine-tuning and mixup pretrain settings
@@
E.5 Ablation study of the mask ratio in referring-aware dynamic masking strategy